OpenStudy (anonymous):

@iGreen help i will meadel!!!!!!! 23/40-9/30 A. 0 b. 1 c. 1/2

OpenStudy (igreen):

Well, you have to find a common denominator..that's 120 here..so we change it to a common denominator of 120. \(\dfrac{23}{40} - \dfrac{9}{30}\) \(\dfrac{23 \times 3}{40 \times 3} - \dfrac{9 \times 4}{30 \times 4}\) \(\dfrac{69}{120} - \dfrac{36}{120}\) Can you subtract those?

OpenStudy (anonymous):

33

OpenStudy (igreen):

Yes..don't forget it's a fraction.

OpenStudy (anonymous):

IS THAT HALF OR 1 OR 0

OpenStudy (igreen):

............

OpenStudy (anonymous):

SO WHAT IS THE ANSWER ON MY QUESTION A B C OR D

OpenStudy (igreen):

I'm not going to give you the answer, you have to figure that out yourself with what I gave you..

OpenStudy (igreen):

69 - 36 is 33..you got that right..but don't forget you keep the bottom of the fraction.

OpenStudy (anonymous):

ALRIGHT

OpenStudy (abhisar):

@oregon8 We can help you - as long as you are looking to learn how to get your solution (and not just the answer).

OpenStudy (igreen):

^